# A faster algorithm for Vertex Cover parameterized by solution size

### v6

@@ 9.2 Basic branching rules
Line 10: Suppose we are branching on a 4-vertex $u$ with neighbors $x_{1},x_{2},y_{1},y_{2}$ where $\deg(x_{1})=5$ and $x_{1}\sim x_{2}$. By Proposition 49, $u$ has no other edges in its neighborhood and $y_{1},y_{2}$ are subquartic. Since $x_{2},y_{1},y_{2}$ are within distance two of 5-vertex $x_{1}$ we have $\deg(y_{1})=\deg(y_{2})=4$ and so by Observation 52 we have $\operatorname{codeg}(x_{1},y_{2})=\operatorname{codeg}(x_{1},y_{1})=1$.
Here, we split on $y_{1}$, and then apply (A2) to $x_{1}$ in subproblem $G-y_{1}$. This generates subproblems $G_{1}=G-N[y_{1}],G_{2}=G-y_{1}-N[y_{2},x_{1}]$, and $G_{3}=G-\{y_{1},x_{1}\}$. Since $G$ is in 5BP, subproblem $G_{1}$ has drop $(1.5,4)$ directly. Also, by Lemma 53, we have $\operatorname{minsurp}(G_{2})\geq\operatorname{shad}(N[y_{2},x_{1}])-1\geq 0$. So $G_{2}$ has drop $(3.5,9)$. Subproblem $G_{3}$ has drop $(1,2)$ directly, and has a 2-vertex $u$ so it has drop $(1,3)$ after simplification. We need:
@@
For Branch5-2, each 5-vertex has at least three 4-neighbors (else we would branch at Line 12), and each 4-vertex has at most one 5-neighbor. So $n_{4}\geq 3n_{5}$, and we need
@@ 9.3 Line 6: A 5-vertex with five independent 5-neighbors
Our basic plan is to choose an ordering of the neighbors of $u$ as $v_{1},\dots,v_{5}$, and then split on $v_{1},v_{2},v_{3}$ to create a 2-vertex $u$. The graph $H=G-\{v_{1},v_{2},v_{3}\}-N[v_{4},v_{5}]$ is central to this analysis; note that $\operatorname{codeg}(v_{4},v_{5})\in\{1,2\}$ by Proposition 47.
Let $I$ be a min-set of $H$, and define parameters
$s=|N(I)|,\qquad\ell=\operatorname{minsurp}(H),\qquad d_{45}=\operatorname{codeg}(v_{4},v_{5})$
For this section, we define $A=\{v_{1},v_{2},v_{3}\}$ and $B=\{v_{4},v_{5}\}$, so that $H=G-(A\cup N[B])$.
@@
If there is any vertex $x\in H$ with $\deg_{H}(x)\leq 5-\deg_{G}(x)$, then $u$ has neighbors $v_{i},v_{j}$ with $R(v_{i},N[v_{j}])\geq 1$.
@@
If $\deg_{G}(x)=3$, then $x\not\sim A$; to get $\deg_{H}(x)\leq 2$, we must have $\operatorname{codeg}(x,v_{j})\geq 1$ for $v_{j}\in B$, so $R(v_{1},N[v_{j}])\geq 1$. If $\deg_{G}(x)=4$, then $|N(x)\cap A|\leq 1$ (else we would branch on $x$ at Line 3); so to get $\deg_{H}(x)\leq 1$, we have either $\operatorname{codeg}(x,v_{j})\geq 2$ for $v_{j}\in B$, or $|N(x)\cap A|\geq 1$. If $\operatorname{codeg}(x,v_{j})\geq 2$ for $v_{j}\in B$, then $R(v_{1},N[v_{j}])\geq 1$ (due to subquadratic vertex $x$), while if $\operatorname{codeg}(x,v_{4})=\operatorname{codeg}(x,v_{5})=1,x\sim v_{i}\in A$, then $R(v_{i},N[v_{4}])\geq 1$ (due to subquadratic vertex $x$).
Finally, suppose $\deg_{G}(x)=5$ and $x$ is isolated in $H$. By Proposition 47 we have $\operatorname{codeg}(x,u)\leq 2$ and $\operatorname{codeg}(x,v_{4})\leq 2$ and $\operatorname{codeg}(x,v_{5})\leq 2$. If $\operatorname{codeg}(x,v_{4})=\operatorname{codeg}(x,v_{5})=2$, then $x$ has a neighbor $v_{i}\in A$, so $R(v_{i},N[v_{4}])\geq 1$ (due to subquadratic vertex $x$). Otherwise, if $\operatorname{codeg}(x,v_{j})=2$ for $v_{j}\in B$ and $x\sim v_{i}\in A$, then $R(v_{i},N[v_{j}])\geq 1$ (due to subquadratic vertex $x$) ∎
@@
  • There are distinct neighbors $v_{i},v_{j}$ where $G-v_{i}-N[v_{j}]$ is 3-vul.
  • $d_{45}=1$ and there are neighbors $v_{i},v_{i^{\prime}},v_{j}$ with $R(v_{i},v_{i^{\prime}},N[v_{j}])\geq 1$.
  • $d_{45}=2$ and the graph $G-N[u]$ has a 5-vertex with a 3-neighbor.
@@
If some $x\in I$ has exactly one neighbor $v_{i}\in A$, then $\operatorname{codeg}(x,v_{j})\geq 2$ for some $v_{j}\in B$ and $R(v_{i},N[v_{j}])\geq 1$ (due to subquadratic vertex $x$).
If $d_{45}=1$, the only remaining case is that there is some $x\in I$ with two neighbors $v_{i},v_{i^{\prime}}$ in $A$. Then $\operatorname{codeg}(x,v_{j})\geq 1$ for some $v_{j}\in B$, and so $R(v_{i},v_{i^{\prime}},N[v_{j}])\geq 1$ as claimed.
Finally suppose $d_{45}=2$. If $|N(I)\cap A|<3$, say without loss of generality $v_{1}\not\sim I$, then $\operatorname{surp}_{G-\{v_{2},v_{3}\}-N[v_{4},v_{5}]}(I)=-1$ and hence $\operatorname{surp}_{G-v_{2}-N[v_{4}]}(I\cup\{v_{5}\})\leq 2$; so $G-v_{2}-N[v_{4}]$ is 3-vul (via indset $I\cup\{v_{5}\})$. So let us suppose that every vertex in $A$ has at least one neighbor in $I$, and no $x\in I$ has exactly one neighbor in $A$. Without loss of generality, we may suppose that $v_{1}\sim x_{1}$ and $v_{2}\sim x_{2}$, and $v_{3}\sim x_{1}$ and $v_{3}\sim x_{2}$, with no other edges between $A$ and $I$.
Now $v_{3}$ has four 5-neighbors $u,x_{1},x_{2},y$ and $y,v_{3}$ share neighbors $x_{1},x_{2}$. So by Proposition 50 we have $y\not\sim v_{3}$. If $y\sim v_{i}$ for $v_{i}\in\{v_{1},v_{2}\}$, then $R(v_{i},N[v_{3}])\geq 1$ due to subquadratic vertex $y$. Otherwise, if $y\not\sim A$, then in the graph $G-N[u]$, vertices $x_{1},x_{2}$ remain with degree 3 (losing neighbors $v_{1},v_{2},v_{3}$), while $y$ remains with degree 5. ∎
We now analyze the subproblems. When describing each case, we assume that no other vertex ordering could go into an earlier case.
Case I: $\boldsymbol{\operatorname{shad}(v_{1},v_{2},N[v_{3}])\leq-1.}$ Let $G^{\prime}=G-\{v_{1},v_{2}\}-N[v_{3}]$. We claim that $\operatorname{mindeg}(G^{\prime})\geq 1$. For, consider a vertex $x\in G^{\prime}$. By Proposition 47, we have $\operatorname{codeg}(x,v_{3})\leq 2$. If $\deg(x)=5$, this implies $\deg_{G^{\prime}}(x)\geq 1$. If $\deg(x)=4$, then $|N(x)\cap\{v_{1},v_{2}\}|\leq 1$ (else we would branch on $x$ at Line 4), so again $\deg_{G^{\prime}}(x)\geq 1$. Finally, if $\deg(x)=3$, then $x\not\sim\{v_{1},v_{2}\}$ and $\deg_{G^{\prime}}(x)\geq 1$.
So $\operatorname{surp}_{G^{\prime}}(J)\leq-1$ for an indset $J$, where necessarily $|J|\geq 2$. Then $\operatorname{surp}_{G-N[v_{3}]}(J)\leq 1$; since $G$ is in 5BP, this implies $R(N[v_{3}])\geq|J|\geq 2$. Our strategy is to split on $v_{3}$, generating subproblems $G-v_{3},G-N[v_{3}]$; these have drops $(0.5,1),(2,7)$ after simplification. We need:
@@
Case II: $\boldsymbol{R(v_{1},v_{2},N[v_{3}])\geq 3.}$ Split on $v_{1},v_{2},v_{3}$, generating subproblems $G-N[v_{1}],G-v_{1}-N[v_{2}],G-\{v_{1},v_{2}\}-N[v_{3}],G-\{v_{1},v_{2},v_{3}\}$. Since we are not in Case I, the first three have drops $(2,5),(2.5,6),(3,7)$ directly. Since $G$ is in 5BP, subproblem $G-\{v_{1},v_{2},v_{3}\}$ has drop $(1.5,3)$ directly; we apply Proposition 30 to its 2-vertex $u$ where $\deg(v_{4})+\deg(v_{5})-d_{45}-1\geq 7$. So we need:
@@
Note: After Case II, we suppose that $\ell\geq-3+d_{45}$. For, suppose $\operatorname{surp}_{H}(I)\leq-4+d_{45}\leq-2$, where necessarily $|I|\geq 2$. Then $\operatorname{surp}_{G-\{v_{1},v_{2}\}-N[v_{4}]}(I\cup\{v_{5}\})\leq 1$. So either $\operatorname{shad}(v_{1},v_{2},N[v_{4}])\leq-1$ (which would put us into Case I with alternate vertex ordering $v_{1},v_{2},v_{4}$) or, by Lemma 13 we have $R(v_{1},v_{2},N[v_{4}])\geq|I\cup\{v_{5}\}|\geq 3$, putting us into Case II.
Our strategy in all subsequent cases is to split on $v_{1},v_{2},v_{3}$, generating subproblems $G_{1}=G-N[v_{1}],G_{2}=G-v_{1}-N[v_{2}],G_{3}=G-\{v_{1},v_{2}\}-N[v_{3}]$, and $G-\{v_{1},v_{2},v_{3}\}$. We then apply (A1) to the 2-vertex $u$ in $G-\{v_{1},v_{2},v_{3}\}$, obtaining subproblems $G_{0}=G-N[u]$ and $H=G-\{v_{1},v_{2},v_{3}\}-N[v_{4},v_{5}]$. Since we are not in Case I, subproblems $G_{0},G_{1},G_{2},G_{3}$ have drops $(2,5),(2,5),(2.5,6)$ and $(3,7)$ respectively, and the final subproblem $H$ has drop
$(5.5-d_{45}/2,13-d_{45})\qquad\text{ or }\qquad(5.5-(d_{45}+\ell)/2,13-d_{45}+s)$
where the latter is obtained if we apply P1 to $I$.
@@
Note: After Case IV, we must have $s\geq 1$; for, if $s=0$, then $H$ has an isolated vertex; by Proposition 55 there are then neighbors $v_{i},v_{j}$ with $R(v_{i},N[v_{j}])\geq 1$, putting us into Case IV.
Case V: $\boldsymbol{G_{3}}$ is 4-vul. We know $\ell\geq-3+d_{45}$. First, suppose $d_{45}=2,\ell=-1,s=1$ exactly. So $H$ has drop $(4,12)$. By Proposition 56, there are either vertices $v_{i},v_{j}$ where $G-v_{i}-N[v_{j}]$ is 3-vul (putting us into Case IV), or $G_{0}$ has a 5-vertex with a 3-neighbor. In the latter case, we can either simplify $G_{0}$ further, or apply Lemma 24 to $G_{0}$, and we then need:
@@
Otherwise, if $\ell\geq-2+d_{45}$ or $s\geq 2$ or $d_{45}=1$, then $H$ has drop $(4.5,11)$ or $(4,13)$, and we need:
@@
Case VI: $\boldsymbol{d_{45}=1}$. If $\ell\leq-2$, then $\operatorname{surp}_{G-\{v_{1},v_{2}\}-N[v_{4}]}(I\cup\{v_{5}\})\leq 2$ and $|I|=s-\ell\geq 3$, so $G-\{v_{1},v_{2}\}-N[v_{4}]$ is 4-vul; this would put us into Case V. If $s=1,\ell=-1$, then by Proposition 56, there are either neighbors $v_{i},v_{i^{\prime}},v_{j}$ with $R(v_{i},v_{i^{\prime}},N[v_{j}])\geq 1$ (putting us into Case V) or $G-v_{i}-N[v_{j}]$ is 3-vul (putting us into Case IV). So either we haver $\ell\geq 0$ or we have $\ell=-1,s\geq 2$, and thus $H$ has drop $(4.5,14)$ or $(5,12)$. We need:

### v7

@@ 7.2 Basic branching rules
Line 10: Let us branch on a 4-vertex $u$ with neighbors $x_{1},x_{2},y_{1},y_{2}$ where $\deg(x_{1})=5$ and $x_{1}\sim x_{2}$. This is Case II of Proposition 50: $u$ has no other edges in its neighborhood and $y_{1},y_{2}$ are subquartic. Since $x_{2},y_{1},y_{2}$ are within distance two of 5-vertex $x_{1}$ we have $\deg(y_{1})=\deg(y_{2})=4$ and so by Observation 53 we have $\operatorname{codeg}(x_{1},y_{2})=\operatorname{codeg}(x_{1},y_{1})=1$.
Here, we split on $y_{1}$, and then apply (B-Fun2) to $x_{1}$ in subproblem $G-y_{1}$. This generates subproblems $G_{1}=G-N[y_{1}],G_{2}=G-y_{1}-N[y_{2},x_{1}]$, and $G_{3}=G-\{y_{1},x_{1}\}$. Subproblem $G_{1}$ has drop $(1.5,4)$. By Lemma 54, $\operatorname{minsurp}(G_{2})\geq\operatorname{shad}(N[y_{2},x_{1}])-1\geq 0$, so $G_{2}$ has drop $(3.5,9)$. Subproblem $G_{3}$ has drop $(1,2)$ directly, and has a 2-vertex $u$ so it has drop $(1,3)$ after simplification. We need:
@@
For Branch5-2, each 5-vertex has at least three 4-neighbors (else we would branch at Line 11), and each 4-vertex has at most one 5-neighbor. So $n_{4}\geq 3n_{5}$, and we need
@@ 7.3 Line 6: A 5-vertex with five independent 5-neighbors
Our basic plan is to choose an ordering of the neighbors of $u$ as $v_{1},\dots,v_{5}$, and then split on $v_{1},v_{2},v_{3}$ to create a 2-vertex $u$. For Section 7.3, we use the following notations:
$A=\{v_{1},v_{2},v_{3}\},\qquad B=\{v_{4},v_{5}\},\qquad d_{45}=\operatorname{codeg}(v_{4},v_{5})$
$H=G-\{v_{1},v_{2},v_{3}\}-N[v_{4},v_{5}]=G-A-N[B]$
$I=\text{a chosen min-set of $H$},\qquad s=|N(I)|,\qquad\ell=\operatorname{minsurp}(H)=\operatorname{surp}_{H}(I)$
The properties of the graph $H$ are central to this analysis; note that $d_{45}\in\{1,2\}$ by Proposition 48. We begin with a few observations.
@@
If there is any vertex $x\in H$ with $\deg_{H}(x)\leq 5-\deg_{G}(x)$, then $u$ has neighbors $v_{i},v_{j}$ with $S(v_{i},N[v_{j}])\geq 1$.
@@
If $\deg_{G}(x)=3$, then $x\not\sim A$; to get $\deg_{H}(x)\leq 2$, we must have $\operatorname{codeg}(x,v_{j})\geq 1$ for $v_{j}\in B$, so $S(v_{1},N[v_{j}])\geq 1$. If $\deg_{G}(x)=4$, then $|N(x)\cap A|\leq 1$ (else we would branch on $x$ at Line 4); so to get $\deg_{H}(x)\leq 1$, we have either $\operatorname{codeg}(x,v_{j})\geq 2$ for $v_{j}\in B$, or $x\sim A$. In the former case, we have $S(v_{1},N[v_{j}])\geq 1$ (due to subquadratic vertex $x$). In the latter case, if $\operatorname{codeg}(x,v_{4})=\operatorname{codeg}(x,v_{5})=1,x\sim v_{i}\in A$, then $S(v_{i},N[v_{4}])\geq 1$ (due to subquadratic vertex $x$).
Finally, suppose $\deg_{G}(x)=5$ and $x$ is isolated in $H$. By Proposition 48 we have $\operatorname{codeg}(x,u)\leq 2$ and $\operatorname{codeg}(x,v_{4})\leq 2$ and $\operatorname{codeg}(x,v_{5})\leq 2$. If $\operatorname{codeg}(x,v_{4})=\operatorname{codeg}(x,v_{5})=2$, then $x$ has a neighbor $v_{i}\in A$, so $S(v_{i},N[v_{4}])\geq 1$ (due to subquadratic vertex $x$). Otherwise, if $\operatorname{codeg}(x,v_{j})=2$ for $v_{j}\in B$ and $x\sim v_{i}\in A$, then $S(v_{i},N[v_{j}])\geq 1$ (due to subquadratic vertex $x$) ∎
@@
  • There are neighbors $v_{i},v_{j}$ where $G-v_{i}-N[v_{j}]$ is 3-vul.
  • $d_{45}=1$ and there are neighbors $v_{i},v_{i^{\prime}},v_{j}$ with $S(v_{i},v_{i^{\prime}},N[v_{j}])\geq 1$.
  • $d_{45}=2$ and the graph GN[u]G-N[u] has a 5-vertex with a 3-neighbor.
@@
If some $x\in I$ has exactly one neighbor $v_{i}\in A$, then $\operatorname{codeg}(x,v_{j})\geq 2$ for some $v_{j}\in B$ and $S(v_{i},N[v_{j}])\geq 1$ (due to subquadratic vertex $x$). Similarly, if there is some $x\in I$ with two neighbors $v_{i},v_{i^{\prime}}$ in $A$, then $\operatorname{codeg}(x,v_{j})\geq 1$ for some $v_{j}\in B$, and so $S(v_{i},v_{i^{\prime}},N[v_{j}])\geq 1$.
The only remaining case if $d_{45}=2$ and every $x\in I$ has at least two neighbors in $A$. If $|N(I)\cap A|<3$, say without loss of generality $v_{1}\not\sim I$, then $\operatorname{surp}_{G-\{v_{2},v_{3}\}-N[v_{4},v_{5}]}(I)=-1$ and hence $\operatorname{surp}_{G-v_{2}-N[v_{4}]}(I\cup\{v_{5}\})\leq 2$; so $G-v_{2}-N[v_{4}]$ is 3-vul (via indset $I\cup\{v_{5}\})$. So we may suppose without loss of generality that the edges between $A$ and $I$ are precisely $(v_{1},x_{1}),(v_{2},x_{2}),(v_{3},x_{1}),(v_{3},x_{2})$.
Now $v_{3}$ has three 5-neighbors $u,x_{1},x_{2}$ and $y,v_{3}$ share neighbors $x_{1},x_{2}$. So by Proposition 51 we have $y\not\sim v_{3}$ (else $v_{3}$ would have four 5-neighbors and its neighborhood would be a non-matching). If $y\sim v_{i}$ for $v_{i}\in\{v_{1},v_{2}\}$, then $S(v_{i},N[v_{3}])\geq 1$ due to subquadratic vertex $y$. Otherwise, if $y\not\sim A$, then in the graph GN[u]G-N[u], vertices $x_{1},x_{2}$ remain with degree 3 (losing neighbors $v_{1},v_{2},v_{3}$), while $y$ remains with degree 5. ∎
We now analyze the branching rule and subproblems. When describing each case, we assume that no other vertex ordering could go into an earlier case.
Case I: $\boldsymbol{\operatorname{shad}(v_{1},v_{2},N[v_{3}])\leq-1.}$ Let $G^{\prime}=G-\{v_{1},v_{2}\}-N[v_{3}]$ and let $J$ be a min-set of $G^{\prime}$. We claim that $\operatorname{mindeg}(G^{\prime})\geq 1$. For, consider a vertex $x\in G^{\prime}$. By Proposition 48, we have $\operatorname{codeg}(x,v_{3})\leq 2$. If $\deg(x)=5$, this implies $\deg_{G^{\prime}}(x)\geq 1$. If $\deg(x)=4$, then $|N(x)\cap\{v_{1},v_{2}\}|\leq 1$ (else we would branch on $x$ at Line 4), so again $\deg_{G^{\prime}}(x)\geq 1$. Finally, if $\deg(x)=3$, then $x\not\sim\{v_{1},v_{2}\}$ and $\deg_{G^{\prime}}(x)\geq 1$.
So necessarily $|J|\geq 2$. Then $\operatorname{surp}_{G-N[v_{3}]}(J)\leq 1$, which implies that $S(N[v_{3}])\geq|J|\geq 2$. Our strategy is to split on $v_{3}$, generating subproblems $G-v_{3},G-N[v_{3}]$ with drops $(0.5,1),(2,7)$ after simplification. We need:
@@
Case II: $\boldsymbol{S(v_{1},v_{2},N[v_{3}])\geq 3.}$ We split on $v_{1},v_{2},v_{3}$, generating subproblems $G-N[v_{1}],G-v_{1}-N[v_{2}],G-\{v_{1},v_{2}\}-N[v_{3}],G-\{v_{1},v_{2},v_{3}\}$. Since we are not in Case I, the first three have drops $(2,5),(2.5,6),(3,7)$ directly. Note that $\operatorname{shad}(v_{1},v_{2},v_{3})\geq\operatorname{shad}(v_{1})-2\geq 0$. So subproblem $G-\{v_{1},v_{2},v_{3}\}$ has drop $(1.5,3)$ directly, and we can apply Proposition 30 to its 2-vertex $u$ where $\deg(v_{4})+\deg(v_{5})-d_{45}-1\geq 7$. So we need:
@@
Note: After Case II, we suppose that $\ell\geq-3+d_{45}$. For, suppose $\operatorname{surp}_{H}(I)\leq-4+d_{45}\leq-2$, where necessarily $|I|\geq 2$. Then $\operatorname{surp}_{G-\{v_{1},v_{2}\}-N[v_{4}]}(I\cup\{v_{5}\})\leq 1$. So either $\operatorname{shad}(v_{1},v_{2},N[v_{4}])\leq-1$ (which would put us into Case I with alternate vertex ordering $v_{1},v_{2},v_{4}$) or, by Lemma 13 we have $S(v_{1},v_{2},N[v_{4}])\geq|I\cup\{v_{5}\}|\geq 3$, putting us into Case II.
Our strategy in all subsequent cases is to split on $v_{1},v_{2},v_{3}$, generating subproblems $G_{1}=G-N[v_{1}],G_{2}=G-v_{1}-N[v_{2}],G_{3}=G-\{v_{1},v_{2}\}-N[v_{3}]$, and $G-\{v_{1},v_{2},v_{3}\}$; we then apply (B-Pair) to the 2-vertex $u$ in $G-\{v_{1},v_{2},v_{3}\}$, to get subproblems $G_{0}=G-N[u]$ and $H=G-\{v_{1},v_{2},v_{3}\}-N[v_{4},v_{5}]$. Since we are not in Case I, subproblems $G_{0},G_{1},G_{2},G_{3}$ have drops $(2,5),(2,5),(2.5,6)$ and $(3,7)$ respectively, and the final subproblem $H$ has drop
$(5.5-d_{45}/2,\ \ \ \ \ \ \ \ 13-d_{45})$ if $\ell\geq 0$
$(5.5-d_{45}/2+\ell/2,13-d_{45}+s)$ if $\ell<0$ (after applying P1)
@@
Note: After Case IV, we must have $s\geq 1$; for, if $s=0$, then $H$ has an isolated vertex; by Proposition 56 there are then neighbors $v_{i},v_{j}$ with $S(v_{i},N[v_{j}])\geq 1$, putting us into Case IV.
Case V: $\boldsymbol{G_{3}}$ is 4-vul. Recall that $\ell\geq-3+d_{45}$. First, suppose $d_{45}=2,\ell=-1,s=1$, so $H$ has drop $(4,12)$. By Proposition 57, there are either vertices $v_{i},v_{j}$ where $G-v_{i}-N[v_{j}]$ is 3-vul (putting us into Case IV), or $G_{0}$ has a 5-vertex with a 3-neighbor. In the latter case, we can either simplify $G_{0}$ further, or apply Lemma 24 to $G_{0}$, and we then need:
@@
Otherwise, if $\ell\geq-2+d_{45}$ or $s\geq 2$ or $d_{45}=1$, then $H$ has drop $(4.5,11)$ or $(4,13)$; we need:
@@
Case VI: $\boldsymbol{d_{45}=1}$. If $\ell\leq-2$, then $\operatorname{surp}_{G-\{v_{1},v_{2}\}-N[v_{4}]}(I\cup\{v_{5}\})\leq 2$ and $|I|=s-\ell\geq 3$, so $G-\{v_{1},v_{2}\}-N[v_{4}]$ is 4-vul; this would put us into Case V. If $s=1,\ell=-1$, then by Proposition 57, there are either neighbors $v_{i},v_{i^{\prime}},v_{j}$ with $S(v_{i},v_{i^{\prime}},N[v_{j}])\geq 1$ (putting us into Case V) or $G-v_{i}-N[v_{j}]$ is 3-vul (putting us into Case IV). So either $\ell\geq 0$ or $\ell=-1,s\geq 2$, and thus $H$ has drop $(5,12)$ or $(4.5,14)$ respectively. We need: